# Rotated Multi-Scale Interaction Network for Referring Remote Sensing Image Segmentation

### v2

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) is a new challenge that combines computer vision and natural language processing, delineating specific regions in aerial images as described by textual queries. Traditional Referring Image Segmentation (RIS) approaches have been impeded by the complex spatial scales and orientations found in aerial imagery, leading to suboptimal segmentation results. To address these challenges, we introduce the Rotated Multi-Scale Interaction Network (RMSIN), an innovative approach designed for the unique demands of RRSIS. RMSIN incorporates an Intra-scale Interaction Module (IIM) to effectively address the fine-grained detail required at multiple scales and a Cross-scale Interaction Module (CIM) for integrating these details coherently across the network. Furthermore, RMSIN employs an Adaptive Rotated Convolution (ARC) to account for the diverse orientations of objects, a novel contribution that significantly enhances segmentation accuracy. To assess the efficacy of RMSIN, we have curated an expansive dataset comprising 17,402 image-caption-mask triplets, which is unparalleled in terms of scale and variety. This dataset not only presents the model with a wide range of spatial and rotational scenarios but also establishes a stringent benchmark for the RRSIS task, ensuring a rigorous evaluation of performance. Our experimental evaluations demonstrate the exceptional performance of RMSIN, surpassing existing state-of-the-art models by a significant margin. All datasets and code are made available at https://github.com/Lsan2401/RMSIN.
@@ 1 Introduction
Referring Remote Sensing Image Segmentation (RRSIS) stands at the forefront of integrating computer vision with natural language processing, aiming to segment specific areas from aerial images based on textual descriptions. This sophisticated task goes beyond the capabilities of traditional RIS [14, 28, 25], requiring an intricate understanding of the spatial and geographic nuances conveyed from aerial perspectives. RRSIS plays a crucial role in a wide range of applications, including land use categorization [9], climate impact studies [38], and urban infrastructure management [8]. By pushing the boundaries of semantic understanding in remote sensing data, RRSIS is advancing the possibilities in these domains. Despite this, the field has been constrained by the limited scale and scope of existing datasets, which are insufficient for training models to the level of accuracy required for these critical tasks.
@@
Furthermore, although the existing RIS methodologies [18, 25, 41, 29] have demonstrated effectiveness in specific domains [53, 33, 35], they face limitations when applied to the diverse and intricate nature of remote sensing imagery. As illustrated in Fig. 1, aerial images pose distinct challenges that are not encountered in conventional image datasets, including vast and diverse spatial scales, as well as objects captured from multiple orientations. Current RIS approaches typically excel in aligning visual and linguistic elements in well-bounded contexts [52, 17] but falter when faced with the chaotic and unstructured nature of aerial images. The inability of these methods to grapple with high levels of spatial variance and rotational diversity results in a notable performance gap in RRSIS tasks, highlighting the need for a more robust and versatile approach.
@@
  • We introduce RRSIS-D, a novel benchmark dataset tailored for Referring Remote Sensing Image Segmentation (RRSIS). This dataset uniquely leverages the powerful segmentation capabilities of the SAM in conjunction with manual calibration, accommodating substantial variations in both spatial resolution and object orientation.
@@
  • Extensive experiments demonstrate the superiority of our proposed RMSIN over current state-of-the-art baselines, consistently achieving enhanced performance across a range of evaluation metrics.
@@ 2 Related work
Referring Image Detection and Segmentation. Referring Image Detection aims to predict a bounding box corresponding to a given referring expression. Existing works can be classified into two-staged methods [13, 15, 56, 58] which are based on region proposal ranking, and one-stage methods [1, 24, 26, 48, 57] which directly predict the target bounding box. Referring Image Segmentation aims to achieve pixel-level localization of target objects within images based on associated referring expressions, presenting a more complex task [14]. Early works [28, 25, 35] leverage convolution networks and recurrent neural networks to extract vision and language features, respectively. These features are then fused by simple concatenation to generate final predictions. Subsequent methods [2, 28, 3, 19, 34, 41, 29, 20, 6] mainly focus on vision-language alignment to enhance predictions. Some employ recurrent refinement [2, 28], while others utilize dynamic filters [3, 19, 34] to fuse visual and linguistic features. Recently, leveraging the Transformer’s outstanding performance, methods can deviled into two categories: those performing cross-modal decoder fusion based on Transformer [41, 29, 20, 6] and those incorporating language-aware visual encoding instead of post-feature fusion [49, 44, 22]. However, due to the specific characteristics of aerial images, these methods exhibit limited performance in the Remote Sensing field. Some approaches [21, 36, 51] have introduced extra scale interaction modules to enhance feature extraction. However, the extreme semantic gap between natural images and aerial images still results in suboptimal performance.
@@ 4.1 Overview
The pipeline of our proposed model is depicted in Fig. 5. Initially, given the input image $I\in\mathbb{R}^{H\times W\times 3}$ and the language expression $E=\{\omega_{i}\},{i\in\{0,\dots,N\}}$, where $H$ and $W$ represent the height and width of the input image, and $N$ is the length of the expression, the input language expression $E$ is transformed into the feature space $F_{\ell}\in\mathbb{R}^{N\times C}$ via the backbone $f_{\ell}$.
The following Compounded Scale Interaction Encoder (CSIE), which is composed of an Intra-scale Interaction Module (IIM) at each stage, and a Cross-scale Interaction Module (CIM), is applied to generate the fused features with sufficient semantics across multiple scales. Finally, we propose an Adaptive Rotated Convolution (ARC) based Oriented-Aware Decoder (OAD) to generate the segmentation mask by the parallel inference on the features from the multiple stages of the CSIE.
@@ 4.2.1 Intra-scale Interaction Module
$\alpha={\rm Tanh(LN(ReLU(LN(}\hat{F}_{e1}^{i-1})))),$ (4)
where ${\rm LN(\cdot)}$ denotes a $1\times 1$ convolution, and ${\rm Tanh}(\cdot)$ and ${\rm ReLU(\cdot)}$ represent the activation functions.
@@
where $W^{i}_{q}$, $W^{i}_{k}$ and $W^{i}_{v}$ are the liner projection matrices. Subsequently, the attention $A^{i}$ is combined with $\hat{F}_{e}^{i-1}$ to obtain language-guided image features:
@@ 4.2.2 Cross-scale Interaction Module
Multi-scale Attention Layer is subsequently implemented. Specifically, we design various perceptive fields for the concatenated feature $F_{c}^{\ast}$ to achieve deep multi-scale interaction. $F_{c}^{\ast}$ is resized to different scales with the assistance of a diverse combination of depth-wise convolution kernel sizes and strides, defined as follows:
$F_{c}^{m}=\mathop{\rm concat}\limits_{c}(k^{m}\ast F_{c}^{\ast})W^{m},$ (9)
@@
where $m\in\{1,\dots,M\}$, $M$ is the number of resized scales, $k_{m}$ is the parameter of the $m$-th depth-wise convolution kernel. $h_{m}$ and $w_{m}$ are the height and weight of the $F_{c}^{m}$. With the set $\{F_{c}^{m}|m\in\{1,\dots,M\}\}$, we flatten all the elements on the size dimension and concatenate them as a sequence features $\hat{F}_{c}^{\ast}\in\mathbb{R}^{(\sum_{1}^{M}h^{m}\times w^{m})\times C}$. Similar to vanilla attention [40], we take the origin feature $F_{c}^{\ast}$ as the query, and the multi-scale-aware feature $\hat{F}_{c}^{\ast}$ as the key and value to perform cross-scale interaction:
@@
where $i\in\{1,2,3,4\}$. The output of the Scale-aware Gate is utilized in the subsequent decoder for final mask prediction.
@@ 4.3.1 Adaptive Rotated Convolution
where ${\rm Seg}(\cdot)$ refers to a nonlinear block comprising a $3\times 3$ convolution layer, a batch normalization layer, and a ReLU activity function to enhance the nonlinearity of the segmentation feature space. And ${\rm Proj}(\cdot)$ is implemented as a liner projection function to map the final feature $D_{1}$ into two class scores. It is notable that half of the convolution layers are replaced by the ARC to leverage orientation information in the feature space, thereby eliminating redundancy for enhanced accuracy in boundary details.
@@ 5 Experiments
Method Visual Encoder Language Encoder RRSIS-D
P@0.5 P@0.6 P@0.7 P@0.8 P@0.9 oIoU mIoU
RRN [25] ResNet-101 [11] LSTM [12] 51.09 42.47 33.04 20.80 6.14 66.53 46.06
CSMA [52] ResNet-101 None 55.68 48.04 38.27 26.55 9.02 69.68 48.85
LSCM [18] ResNet-101 LSTM 57.12 48.04 37.87 26.37 7.93 69.28 50.36
CMPC [17] ResNet-101 LSTM 57.93 48.85 38.50 25.28 9.31 70.15 50.41
BRINet [16] ResNet-101 LSTM 58.79 49.54 39.65 28.21 9.19 70.73 51.14
CMPC$+$ [30] ResNet-101 LSTM 59.19 49.36 38.67 25.91 8.16 70.14 51.41
LGCE [54] Swin-B [31] BERT [5] 65.23 58.05 49.25 39.77 23.05 74.20 57.30
LAVT [49] Swin-B BERT 69.54 63.51 53.16 43.97 24.25 77.59 61.46
RMSIN (Ours) Swin-B BERT 74.66 68.22 57.41 45.29 24.43 78.27 65.10
Table 2: Comparison with state-of-the-art methods on the validation of our dataset. The best result is bold.
@@
Options P@0.5 P@0.7 P@0.9 oIoU mIoU
Default 69.54 53.16 24.25 77.59 61.46
$+$ Multi-scale Attention 68.91 53.68 25.11 77.61 61.46
$+$ Feed Forward 69.83 52.70 25.57 77.85 61.46
$+$ LRC 73.68 56.67 25.69 77.40 64.25
5.1 Implementation Details
Experiment Settings. In our experiments, the visual backbone utilizes Swin Transformer [31], pre-trained on ImageNet22K [4], while the language backbone employs the base BERT model from HuggingFace’s library [42]. The model is trained for 40 epochs using AdamW [32] with a weight decay of 0.01 and a starting learning rate of 5e-4, reducing according to polynomial decay. Images are resized to $480\times 480$.
Metrics. We follow established protocols by using Overall Intersection-over-Union (oIoU), Mean Intersection-over-Union (mIoU), and Precision@X (P@X) as evaluation metrics, similar to prior studies [55, 43]. oIoU measures the overall ratio of intersection to union areas between predicted and ground truth masks. mIoU calculates average accuracy across all image-region pairs, ensuring equitable assessment of objects regardless of size. P@X evaluates the precision at specific IoU thresholds, reflecting the method’s accuracy in object targeting.
5.2 Comparison with state-of-the-art RIS methods
In our experiments, we compared RMSIN’s performance with existing state-of-the-art referring image segmentation methods for natural images on our RRSIS-D dataset (see Tab. 2). We adhered to the original implementation details of these methods for an equitable comparison. RMSIN outperforms all compared methods across every metric on the validation subset. Notably, it improves mIoU by 3.54% over the recent top-performing method LAVT. This significant enhancement is particularly evident in handling complex cases, such as very small or rotated objects, as shown by gains of 5.12%, 4.71%, and 4.25% in P@0.5, P@0.6, and P@0.7, respectively. These results highlight RMSIN’s ability to capture detailed local and orientation-specific information, leading to more accurate segmentation.
@@ 5.3 Ablation study
We have performed various ablation experiments on RRSIS-D to assess the efficacy of the pivotal components within our proposed network.
Effictiveness of IIM and CIM. To validate the efficacy of our proposed two-scale interaction modules in CSIE, we conduct ablation studies on all the combinations of IIM and CIM. As illustrated in Tab. 4, The introduction of the IIM brings about discernible improvements in precision at lower IoU thresholds, while the incorporation of the CIM further refines predictions across various IoU levels. The combined effect of both modules demonstrates a synergistic enhancement, yielding the highest performance across all evaluated metrics, particularly in P@0.5, P@0.7, and mIoU, with margins ranging from 3.5% to 4.5%. These findings affirm the pivotal role played by the IIM and CIM in capturing multi-scale features from images, thus substantiating their efficacy in advancing the overall segmentation capabilities of the model.
@@
Design options of ARC. We further investigate the impact of the Adaptive Rotated Convolution (ARC) replacement strategy on the results, as demonstrated in Tab. 6 (a). We progressively replace the convolution layers in each stage of the decoder, and the result exhibits a consistent upward trend. Consequently, we opt to replace all three layers of the decoder. Additionally, we explore the influence of varying the number of prediction angles for ARC on the prediction results illustrated in Tab. 6 (b). The decoder showcases its optimal performance when the predicted number of angles is set to 4, resulting in a performance boost of approximately 1% compared to using 1 angle, thereby establishing it as the default setting.

### v3

@@ Abstract
Referring Remote Sensing Image Segmentation (RRSIS) is a new challenge that combines computer vision and natural language processing. Traditional Referring Image Segmentation (RIS) approaches have been impeded by the complex spatial scales and orientations found in aerial imagery, leading to suboptimal segmentation results. To address these challenges, we introduce the Rotated Multi-Scale Interaction Network (RMSIN), an innovative approach designed for the unique demands of RRSIS. RMSIN incorporates an Intra-scale Interaction Module (IIM) to effectively address the fine-grained detail required at multiple scales and a Cross-scale Interaction Module (CIM) for integrating these details coherently across the network. Furthermore, RMSIN employs an Adaptive Rotated Convolution (ARC) to account for the diverse orientations of objects, a novel contribution that significantly enhances segmentation accuracy. To assess the efficacy of RMSIN, we have curated an expansive dataset comprising 17,402 image-caption-mask triplets, which is unparalleled in terms of scale and variety. This dataset not only presents the model with a wide range of spatial and rotational scenarios but also establishes a stringent benchmark for the RRSIS task, ensuring a rigorous evaluation of performance. Experimental evaluations demonstrate the exceptional performance of RMSIN, surpassing existing state-of-the-art models by a significant margin. Datasets and code are available at https://github.com/Lsan2401/RMSIN.
@@ 1 Introduction
Referring Remote Sensing Image Segmentation (RRSIS) stands at the forefront of integrating computer vision with natural language processing [37, 38, 18, 36], aiming to segment specific areas from aerial images based on textual descriptions. This sophisticated task goes beyond the capabilities of traditional RIS [13, 28, 25], requiring an intricate understanding of the spatial and geographic nuances conveyed from aerial perspectives. RRSIS plays a crucial role in a wide range of applications, including land use categorization [9], climate impact studies [44], and urban infrastructure management [8]. By pushing the boundaries of semantic understanding in remote sensing data, RRSIS is advancing the possibilities in these domains. Despite this, the field has been constrained by the limited scale and scope of existing datasets, which are insufficient for training models to the level of accuracy required for these critical tasks.
@@
Furthermore, although the existing RIS methodologies [17, 25, 48, 29] have demonstrated effectiveness in specific domains [62, 39, 41], they face limitations when applied to the diverse and intricate nature of remote sensing imagery. As illustrated in Fig. 1, aerial images pose distinct challenges that are not encountered in conventional image datasets, including vast and diverse spatial scales, as well as objects captured from multiple orientations. Current RIS approaches typically excel in aligning visual and linguistic elements in well-bounded contexts [61, 16] but falter when faced with the chaotic and unstructured nature of aerial images. The inability of these methods to grapple with high levels of spatial variation and rotational diversity results in a notable performance gap in RRSIS tasks, highlighting the need for a more robust and versatile approach.
@@
  • We introduce RRSIS-D, a novel benchmark dataset tailored for Referring Remote Sensing Image Segmentation (RRSIS). This dataset accommodates substantial variations in both spatial resolution and object orientation.
@@ 2 Related work
Referring Image Detection and Segmentation. Referring Image Detection aims to predict a bounding box corresponding to a given referring expression. Existing works can be classified into two-staged methods [12, 14, 65, 67] which are based on region proposal ranking, and one-stage methods [1, 24, 26, 57, 33, 34, 35, 66] which directly predict the target bounding box. Referring Image Segmentation aims to achieve pixel-level localization of target objects within images based on associated referring expressions, presenting a more complex task [13]. Early works [28, 25, 41] leverage convolution networks and recurrent neural networks to extract vision and language features, respectively. These features are then fused by simple concatenation to generate final predictions. Subsequent methods [2, 28, 3, 19, 40, 48, 29, 20, 6, 47, 51, 54] mainly focus on vision-language alignment to enhance predictions. Some employ recurrent refinement [2, 28], while others utilize dynamic filters [3, 19, 40] to fuse visual and linguistic features. Recently, leveraging the Transformer’s outstanding performance, methods can deviled into two categories: those performing cross-modal decoder fusion based on Transformer [48, 29, 20, 6] and those incorporating language-aware visual encoding instead of post-feature fusion [58, 52, 22]. However, due to the specific characteristics of aerial images, these methods exhibit limited performance in the Remote Sensing field. Some approaches [21, 42, 60] have introduced extra scale interaction modules to enhance feature extraction. However, the extreme semantic gap between natural images and aerial images still results in suboptimal performance.
@@ 4.1 Overview
The pipeline of our proposed model is depicted in Fig. 5. Initially, given the input image $I\in\mathbb{R}^{H\times W\times 3}$ and the language expression $E=\{\omega_{i}\},{i\in\{0,\dots,N\}}$, where $H$ and $W$ represent the height and width of the input image, and $N$ is the length of the expression, the input language expression $E$ is transformed into the feature space $F_{\ell}\in\mathbb{R}^{N\times C}$ via the backbone $f_{\ell}$. The following Compounded Scale Interaction Encoder (CSIE), which is composed of an Intra-scale Interaction Module (IIM) at each stage, and a Cross-scale Interaction Module (CIM), is applied to generate the fused features with sufficient semantics across multiple scales. Finally, we propose an Adaptive Rotated Convolution (ARC) based Oriented-Aware Decoder (OAD) to generate the segmentation mask by the parallel inference on the features from the multiple stages of the CSIE.
@@ 4.2.1 Intra-scale Interaction Module
$\alpha={\rm Tanh(Linear(ReLU(Linear(}\hat{F}_{e1}^{i-1})))),$ (4)
where ${\rm Linear(\cdot)}$ denotes the linear projection, and ${\rm Tanh}(\cdot)$ and ${\rm ReLU(\cdot)}$ represent the activation functions.
@@
where $W^{i}_{q}$, $W^{i}_{k}$ and $W^{i}_{v}$ are the linear projection matrices. Subsequently, the attention $A^{i}$ is combined with $\hat{F}_{e}^{i-1}$ to obtain language-guided image features:
@@ 4.2.2 Cross-scale Interaction Module
Multi-scale Attention Layer is subsequently implemented. Specifically, we design various perceptive fields for the concatenated feature $F_{c}^{\ast}$ to achieve deep multi-scale interaction. $F_{c}^{\ast}$ is resized to different scales through the depth-wise convolutions with diverse kernel sizes and strides, defined as follows:
$F_{c}^{m}=k^{m}\ast F_{c}^{\ast},$ (9)
@@
where $m\in\{1,\dots,M\}$, $M$ is the number of resized scales, $k_{m}$ is the $m$-th depth-wise convolution and $h_{m}$ and $w_{m}$ are the corresponding height and weight of the $F_{c}^{m}$. With the set $\{F_{c}^{m}|m\in\{1,\dots,M\}\}$, we flatten all the elements on the size dimension and concatenate them as a sequence features $\hat{F}_{c}^{\ast}\in\mathbb{R}^{(\sum_{1}^{M}h^{m}\times w^{m})\times C}$. Similar to vanilla attention [46], we take the origin feature $F_{c}^{\ast}$ as the query, and the multi-scale-aware feature $\hat{F}_{c}^{\ast}$ as the key and value to perform cross-scale interaction:
@@
where $i\in\{1,2,3,4\}$. The output of the Scale-aware Gate is utilized in the decoder for final mask prediction.
@@ 4.3.1 Adaptive Rotated Convolution
where ${\rm Seg}(\cdot)$ refers to a nonlinear block comprising a $3\times 3$ convolution layer, a batch normalization layer, and a ReLU activity function to enhance the nonlinearity of the segmentation feature space. And ${\rm Proj}(\cdot)$ is implemented as a linear projection function to map the final feature $D_{1}$ into two class scores. It is notable that half of the convolution layers are replaced by the ARC to leverage orientation information in the feature space, thereby eliminating redundancy for enhanced accuracy in boundary details.
@@ 5 Experiments
Method Visual Encoder Text Encoder P@0.5 P@0.6 P@0.7 P@0.8 P@0.9 oIoU mIoU
Val Test Val Test Val Test Val Test Val Test Val Test Val Test
RRN [25] R-101 LSTM 51.09 51.07 42.47 42.11 33.04 32.77 20.80 21.57 6.14 6.37 66.53 66.43 46.06 45.64
CSMA [61] R-101 None 55.68 55.32 48.04 46.45 38.27 37.43 26.55 25.39 9.02 8.15 69.68 69.39 48.85 48.54
LSCM [17] R-101 LSTM 57.12 56.02 48.04 46.25 37.87 37.70 26.37 25.28 7.93 8.27 69.28 69.05 50.36 49.92
CMPC [16] R-101 LSTM 57.93 55.83 48.85 47.40 38.50 36.94 25.28 25.45 9.31 9.19 70.15 69.22 50.41 49.24
BRINet [15] R-101 LSTM 58.79 56.90 49.54 48.77 39.65 39.12 28.21 27.03 9.19 8.73 70.73 69.88 51.14 49.65
CMPC$+$ [30] R-101 LSTM 59.19 57.65 49.36 47.51 38.67 36.97 25.91 24.33 8.16 7.78 70.14 68.64 51.41 50.24
LGCE [63] Swin-B BERT 68.10 67.65 60.52 61.53 52.24 51.45 42.24 39.62 23.85 23.33 76.68 76.34 60.16 59.37
LAVT [58] Swin-B BERT 69.54 69.52 63.51 63.63 53.16 53.29 43.97 41.60 24.25 24.94 77.59 77.19 61.46 61.04
RMSIN (Ours) Swin-B BERT 74.66 74.26 68.22 67.25 57.41 55.93 45.29 42.55 24.43 24.53 78.27 77.79 65.10 64.20
Table 2: Comparison with state-of-the-art methods on the proposed RRSIS-D dataset. R-101 and Swin-B represent ResNet-101 [11] and base Swin Transfomer [31] models, respectively. The best result is bold.
5.1 Implementation Details
Experiment Settings. In our experiments, the visual backbone utilizes Swin Transformer [31], pre-trained on ImageNet22K [4], while the language backbone employs the base BERT model from HuggingFace’s library [49]. The model is trained for 40 epochs using AdamW [32] with a weight decay of 0.01 and a starting learning rate of 3e-5, reducing according to polynomial decay. The setup ran on four RTX 2080 GPUs with a batch size of 8.
Metrics. We utilize Overall Intersection-over-Union (oIoU), Mean Intersection-over-Union (mIoU), and Precision@X (P@X) as evaluation metrics, similar to prior studies [64, 50].
5.2 Comparison with state-of-the-art RIS methods
In our experiments, we compared RMSIN’s performance with existing state-of-the-art referring image segmentation methods on the validation and test subsets of our RRSIS-D dataset (see Tab. 2). For a fair comparison, we adopted the original implementation details of these competing methods. Notably, RMSIN outperforms its counterparts across almost all metrics on both subsets, marking a significant improvement with a 3.64% and 3.16% increase in mIoU on the validation and test subsets respectively over the closest competitor, LAVT. This leap in performance is particularly evident in complex scenarios, such as detecting small or rotated objects, where it secured over 3.0% gains in Precision at IoU thresholds of 0.5, 0.6, and 0.7.
@@ 5.3 Ablation study
We have performed various ablation experiments on the validation subset of RRSIS-D to assess the efficacy of the pivotal components within our proposed network.
Effictiveness of IIM and CIM. To validate the efficacy of our proposed two-scale interaction modules in CSIE, we conduct ablation studies on all the combinations of IIM and CIM. As illustrated in Tab. 3, The introduction of the IIM brings about discernible improvements in precision at lower IoU thresholds, while the incorporation of the CIM further refines predictions across various IoU levels. The combined effect of both modules demonstrates a synergistic enhancement, yielding the highest performance across all evaluated metrics, particularly in P@0.5, P@0.7, and mIoU, with margins ranging from 3.5% to 4.5%. These findings affirm the pivotal role played by the IIM and CIM in capturing multi-scale features from images, thus substantiating their efficacy in advancing the overall segmentation capabilities.
@@
Options P@0.5 P@0.7 P@0.9 oIoU mIoU
Default 69.54 53.16 24.25 77.59 61.46
$+$ Multi-scale Attention 68.91 53.68 25.11 77.61 61.46
$+$ Feed Forward 69.83 52.70 25.57 77.85 61.46
$+$ LRC 73.68 56.67 25.69 77.40 64.25
Design options of ARC. We further investigate the impact of the Adaptive Rotated Convolution (ARC) replacement strategy on the results, as demonstrated in Tab. 6 (a). We progressively replace the convolution layers in each stage of the decoder, and the result exhibits a consistent upward trend. Consequently, we opt to replace all three layers of the decoder. Additionally, we explore the influence of varying the number of prediction angles for ARC on the prediction results illustrated in Tab. 6 (b). The decoder showcases a consistent improvement in performance with an increase in the predicted number of angles, resulting in a performance boost of approximately 1% when seated to 4 compared to 1.